# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

### v1

@@ E.1 Experimental Setup
For text quality, we employed different metrics for different datasets. For the C4 dataset, we utilized perplexity (PPL) and coherence [9] to measure the text quality. For HumanEval, we employed PPL and pass@$k$ score of the code [4]. The pass@$k$ score measures the normalized percentage of solved problems in HumanEval. Formally, the pass score is calculated as
@@
Baselines. We compared against biased and unbiased watermarks in terms of text quality and watermark strength. For further details of baselines, we refer readers to Appendix A. We implemented all LLMs with the Hugging Face library [32]. All watermark benchmarks including KGW, $\gamma$-reweight, and Dipmark were implemented using their public codes.
Implementation details. For all baselines and our methods, we utilized multinomial sampling during text generation. For C4, we employed LLaMA-2-7B as our generative LLM [29]. Following previous work [12], we continued to sample prompts from C4 until we had generated at least 500 text sequences, each consisting of $T=200\pm 5$ tokens. We leveraged LLaMA-2-13B to compute the perplexity of the generated texts. For HumanEval, we applied CodeLLaMA-7B-Instruct [26] as the generative LLM to generate codes for all Python problems. We also leveraged LLaMA-2-13B to compute the perplexity. All experiments were conducted on a single Nvidia A100 GPU with 80GB memory.
@@
[]delimited-[][$\dots$][ … ] Thomas will be responsible for overseeing Micron’s solid state storage business that ranges from hard disk drive replacements with solid state drives (SSDs) to enterprise-class storage solutions. He brings more than 30 years of experience to Micron and most recently served as the vice president of Enterprise Storage for Micron’s common stock is traded on the NASDAQ under the MU symbol. To learn more about Micron Technology, Inc., visit www.micron.com. Micron and the Micron orbit logo are trademarks of Micron Technology, Inc. All other trademarks are the property of their respective owners. []delimited-[][$\dots$][ … ] (PPL:3.25) the Americas region for Seagate Technology. He is a senior executive level leader with a proven track record in defining strategy that drives revenue, profit and new technology execution. "Micron is thrilled to have Darren as part of our team," said Mary Jane Raymond, . []delimited-[][$\dots$][ … ] (PPL:4.45) Fusion I/O, LLC. Before that, Thomas was at Western Digital Corporation where he was a progressive executive, holding various management roles since 2008, most recently as its executive vice president of storage technology. []delimited-[][$\dots$][ … ]
@@
from typing import List def filter_by_prefix (strings: List[str], prefix: str) -> List[str]: """ Filter an input list of strings only for ones that start with a given prefix. $>$ filter_by_prefix([], ’a’) [] $>$ filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’abc’, ’array’] """ return [x for x in strings if x.startswith(prefix)] (Pass, PPL:1.92) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -> List[str]: """ Filter an input list of strings only for ones that start with a given prefix. $>$ filter_by_prefix([], ’a’) [] $>$ filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’array’, ’abc’] """ return [word for word in strings if word.startswith(prefix)] (Pass, PPL:2.18) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -> List[str]: """* Filter an input list of strings only for ones that start with a given prefix. Input: ‘strings = ["abc", "bcd", "cde", "array"]‘, ‘prefix = ’a’‘ Output: ‘["abc", "array"]‘ """ return [word for word in strings if word.startswith(prefix)]
@@
from typing import List def intersperse (numbers: List[int], delimeter: int) -> List[int]: """ Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$ intersperse([], 4) [] $>$ intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] """ if not numbers: return [] result = [] for n in numbers[:-1]: result.append(n) result.append(delimeter) result.append(numbers[-1]) return result (Not pass, PPL:2.45) def intersperse (numbers: List[int], delimeter: int) -> List[int]: """ Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$ intersperse([], 4) [] $>$ intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] """ answer=[] for i in range(0,len(numbers),1): answer = answer + [ numbers[i] ] answer = answer + [ delimeter ] return answer (Not pass, PPL:3.38) Maybe you meant () -> List[int]
@@
def is_prime(n): """Return true if a given number is prime, and false otherwise. ”’ if n < 2: return False for k in range(2, n - 1): if n % k == 0: return False return True (Pass, PPL:1.94) “‘python def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True (Pass, PPL:1.78) def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True if __name__ == ’__main__’: import doctest doctest.testmod()
E.2 Robustness Check on Entropy Threshold Parameter
@@
In this section, we conducted a robustness check on the parameter $\tau$ in STA-M. In particular, we set the low entropy threshold $\tau$ from 0 to 2.1 with an interval of 0.15. At each generation step, we sampled at most 4, 8, and 16 times (i.e., STA-4, STA-8, and STA-16) when the entropy was above the threshold $\tau$. Figure 2 shows text quality and watermark strength of STA-M with different $\tau$s. As depicted, different $\tau$s do not affect the watermark strength significantly for C4 because C4 is a high-entropy dataset. Also, we observe a decrease in PPL when we increase $\tau$ in Figure 2(a), 2(b), and 2(c). The reason is that by setting up a higher entropy threshold, fewer generation steps will apply the STA-M strategy, making the watermarking method more similar to STA-1. According to Figure 2(d), 2(e), and 2(f), we observe a general increase of watermark strength if we have a larger $\tau$ because we will have more green list tokens if we sample M times instead of once. However, higher watermark strength leads to a lower pass@$1$ score, which is related to the text quality [12]. We chose the Pareto optimal of each dataset as our final parameter for each dataset. Specifically, we selected $\tau=1.35$ for C4 and $\tau=1.95$ for HumanEval.
E.3 Examples of STA-generated Texts
@@ E.4 Attacking Watermarks
We introduce the implementation of different attacks as follows. For the GPT-3.5 attack, we utilized the prompt ‘Rewrite the following paragraph: [INPUT]’ for GPT-3.5. For DIPPER-1 [14], we set the lexical diversity to 60 without considering order diversity. Additionally, we increased the order diversity by 20 for DIPPER-2 following previous work [17]. For the copy-paste attack, we randomly replaced $25\%$ of tokens in the watermarked text with tokens from non-watermarked text generated from the same prompt [12].
Appendix F Related Work
@@
Existing white-box watermarking techniques fall into two categories: watermarking during logits and probabilities generation, and watermarking by sampling candidate tokens.
@@
Watermarking during logits and probabilities generation. This category of watermarking methods inserts watermarks into LLMs by artificially adjusting the raw logits or probabilities generated by the LLM. Among this category, Kirchenbauer et al. [12] propose the first watermarking method based on logits adjustment. Their approach randomly partitions the vocabulary set into a green and a red list at each generation step, increasing the logits of green list tokens while keeping red list tokens’ logits fixed. Lee et al. [16] extend the green and red list-based watermarking method to low-entropy scenarios. They adjust the logits only during high-entropy generation steps, leaving the raw logits unchanged for low-entropy steps. Ren et al. [25] improve the vocabulary set partition process by determining the green and red lists based on semantic embeddings of preceding tokens rather than their hash values. Fernandez et al. [8] propose a multi-bit watermarking method that generates a multi-dimensional vector at each generation step, which is utilized to modify logits produced by the original LLM. Their approach allows embedding any bit of watermarking information, up to the dimension of the vector used in the logits adjustment. Yoo et al. [34] develop a multi-bit method by extending the two-list partition idea to multi-list partitions. At each generation step, the vocabulary set is divided into multiple lists. Based on the message to be inserted, the logits for tokens in a selected list are increased, while the token logits in all other lists remain unchanged.
@@
Instead of splitting the vocabulary set into different lists, Hu et al. [11] introduce a method that randomly shuffles the order of all token probabilities within the interval $[0,1]$, setting the probabilities in the first half of the interval to 0 and doubling those in the second half. During the detection phase, a likelihood ratio test examines the significance of the likelihood that the given text is generated with the adjusted probability distribution. Wu et al. [33] further generalizes this method by introducing a hyperparameter $\alpha\in[0,0.5]$, which controls the two cutoff points $\alpha$ and $1-\alpha$ within the interval $[0,1]$. The probability masses for the three resulting sub-intervals are adjusted accordingly.
@@
Watermarking by sampling candidate tokens. This category of watermarking methods inserts watermarks into the token sampling process by using watermark information to control the sampling of candidate tokens. For example, Christ et al. [5] introduce a watermarking method that represents each token in the vocabulary set as a binary string of 0s and 1s. Next, a sequence of values from 0 to 1 is sampled uniformly. These values guide the token sampling process: if the predicted probability for a position in the binary string is larger than the corresponding pseudo-random value, that position is assigned a 1; otherwise, it is assigned a 0. Once all positions are determined, the token corresponding to the resulting binary string is sampled. Additionally, previous work [15] use a sequence of values randomly sampled from a uniform distribution between 0 and 1. The value controls the token sampling process through a decoder function, where the decoder function varies based on the sampling strategy. Hou et al. [10] sample new sentences according to the original LLM until a sentence’s semantic value falls into the acceptance region. The acceptance region is predefined by randomly splitting the space of semantic embedding according to the context and the key.

### v2

@@ F.1 Experimental Setup
For text quality, we employed different metrics for different datasets. For the C4 dataset, we utilized perplexity (PPL) and coherence (Gao et al., 2021) to measure the text quality. For HumanEval, we employed PPL and pass@$k$ score of the code (Chen et al., 2021). The pass@$k$ score measures the normalized percentage of solved problems in HumanEval. Formally, the pass score is calculated as
@@
Baselines. We compared against biased and unbiased watermarks in terms of text quality and watermark strength. For further details of baselines, we refer readers to Appendix B. We implemented all LLMs with the Hugging Face library (Wolf et al., 2019). All watermark benchmarks including KGW, RDW, $\gamma$-reweight, and Dipmark were implemented using their public codes.
Implementation details. For all baselines and our methods, we utilized multinomial sampling during text generation. For C4, we employed LLaMA-2-7B as our generative LLM (Touvron et al., 2023). Following previous work (Kirchenbauer et al., 2023a), we continued to sample prompts from C4 until we had generated at least 500 text sequences, each consisting of $T=200\pm 5$ tokens. We leveraged LLaMA-2-13B to compute the perplexity of the generated texts. For HumanEval, we applied CodeLLaMA-7B-Instruct (Roziere et al., 2023) as the generative LLM to generate codes for all Python problems. We also leveraged LLaMA-2-13B to compute the perplexity. All experiments were conducted on a single Nvidia A100 GPU with 80GB memory.
@@
[]delimited-[][$\dots$][ … ] Thomas will be responsible for overseeing Micron’s solid state storage business that ranges from hard disk drive replacements with solid state drives (SSDs) to enterprise-class storage solutions. He brings more than 30 years of experience to Micron and most recently served as the vice president of Enterprise Storage for Micron’s common stock is traded on the NASDAQ under the MU symbol. To learn more about Micron Technology, Inc., visit www.micron.com. Micron and the Micron orbit logo are trademarks of Micron Technology, Inc. All other trademarks are the property of their respective owners. []delimited-[][$\dots$][ … ] (PPL:3.25) the Americas region for Seagate Technology. He is a senior executive level leader with a proven track record in defining strategy that drives revenue, profit and new technology execution. ”Micron is thrilled to have Darren as part of our team,” said Mary Jane Raymond, . []delimited-[][$\dots$][ … ] (PPL:4.45) Fusion I/O, LLC. Before that, Thomas was at Western Digital Corporation where he was a progressive executive, holding various management roles since 2008, most recently as its executive vice president of storage technology. []delimited-[][$\dots$][ … ]
@@
from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: ””” Filter an input list of strings only for ones that start with a given prefix. $>$much-greater-thanabsent>>>> > > filter_by_prefix([], ’a’) [] $>$much-greater-thanabsent>>>> > > filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’abc’, ’array’] ””” return [x for x in strings if x.startswith(prefix)] (Pass, PPL:1.92) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: ””” Filter an input list of strings only for ones that start with a given prefix. $>$much-greater-thanabsent>>>> > > filter_by_prefix([], ’a’) [] $>$much-greater-thanabsent>>>> > > filter_by_prefix([’abc’, ’bcd’, ’cde’, ’array’], ’a’) [’array’, ’abc’] ””” return [word for word in strings if word.startswith(prefix)] (Pass, PPL:2.18) from typing import List def filter_by_prefix (strings: List[str], prefix: str) -$>$ List[str]: ”””* Filter an input list of strings only for ones that start with a given prefix. Input: ‘strings = [”abc”, ”bcd”, ”cde”, ”array”]‘, ‘prefix = ’a’‘ Output: ‘[”abc”, ”array”]‘ ””” return [word for word in strings if word.startswith(prefix)]
@@
from typing import List def intersperse (numbers: List[int], delimeter: int) -$>$ List[int]: ””” Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$much-greater-thanabsent>>>> > > intersperse([], 4) [] $>$much-greater-thanabsent>>>> > > intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] ””” if not numbers: return [] result = [] for n in numbers[:-1]: result.append(n) result.append(delimeter) result.append(numbers[-1]) return result (Not pass, PPL:2.45) def intersperse (numbers: List[int], delimeter: int) -$>$ List[int]: ””” Insert a number ’delimeter’ between every two consecutive elements of input list ‘numbers’ $>$much-greater-thanabsent>>>> > > intersperse([], 4) [] $>$much-greater-thanabsent>>>> > > intersperse([1, 2, 3], 4) [1, 4, 2, 4, 3] ””” answer=[] for i in range(0,len(numbers),1): answer = answer + [ numbers[i] ] answer = answer + [ delimeter ] return answer (Not pass, PPL:3.38) Maybe you meant () -$>$ List[int]
@@
def is_prime(n): ”””Return true if a given number is prime, and false otherwise. ”’ if n ¡ 2: return False for k in range(2, n - 1): if n % k == 0: return False return True (Pass, PPL:1.94) “‘python def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True (Pass, PPL:1.78) def is_prime(n): if n == 1: return False for i in range(2, n): if n % i == 0: return False return True if __name__ == ’__main__’: import doctest doctest.testmod()
F.2 Robustness Check on Entropy Threshold Parameter
@@
In this section, we conducted a robustness check on the parameter $\tau$ in STA-M. In particular, we set the low entropy threshold $\tau$ from 0 to 2.1 with an interval of 0.15. At each generation step, we sampled at most 4, 8, and 16 times (i.e., STA-4, STA-8, and STA-16) when the entropy was above the threshold $\tau$. Figure 1 shows text quality and watermark strength of STA-M with different $\tau$s. As depicted, different $\tau$s do not affect the watermark strength significantly for C4 because C4 is a high-entropy dataset. Also, we observe a decrease in PPL when we increase $\tau$ in Figure 1(a), 1(b), and 1(c). The reason is that by setting up a higher entropy threshold, fewer generation steps will apply the STA-M strategy, making the watermarking method more similar to STA-1. According to Figure 1(d), 1(e), and 1(f), we observe a general increase of watermark strength if we have a larger $\tau$ because we will have more green list tokens if we sample M times instead of once. However, higher watermark strength leads to a lower pass@$1$ score, which is related to the text quality (Kirchenbauer et al., 2023a). We chose the Pareto optimal of each dataset as our final parameter for each dataset. Specifically, we selected $\tau=1.35$ for C4 and $\tau=1.95$ for HumanEval.
F.3 Examples of STA-generated Texts
@@ F.4 Attacking Watermarks
We introduce the implementation of different attacks as follows. For the copy-paste attack, we randomly replaced $25\%$ of tokens in the watermarked text with tokens from non-watermarked text generated from the same prompt (Kirchenbauer et al., 2023a). For the GPT-3.5 attack, we utilized the prompt ‘Rewrite the following paragraph: [INPUT]’ for GPT-3.5. For DIPPER-1 (Krishna et al., 2024), we set the lexical diversity to 60 without considering order diversity. Additionally, we increased the order diversity by 20 for DIPPER-2 following previous work (Liu et al., 2023a).
Appendix G Related Work
@@
Existing white-box watermarking techniques fall into two categories: watermarking during logits and probabilities generation, and watermarking by controlling sampling strategies.
@@
Watermarking during logits and probabilities generation. This category of watermarking methods inserts watermarks into LLMs by artificially adjusting the raw logits or probabilities generated by the LLM. Among this category, Kirchenbauer et al. (2023a) propose the first watermarking method based on logits adjustment. Their approach randomly partitions the vocabulary set into a green and a red list at each generation step, increasing the logits of green list tokens while keeping red list tokens’ logits fixed. Lee et al. (2023) extend the green and red list-based watermarking method to low-entropy scenarios. They adjust the logits only during high-entropy generation steps, leaving the raw logits unchanged for low-entropy steps. Ren et al. (2023) improve the vocabulary set partition process by determining the green and red lists based on semantic embeddings of preceding tokens rather than their hash values. Fernandez et al. (2023) propose a multi-bit watermarking method that generates a multi-dimensional vector at each generation step, which is utilized to modify logits produced by the original LLM. Their approach allows embedding any bit of watermarking information, up to the dimension of the vector used in the logits adjustment. Yoo et al. (2023) develop a multi-bit method by extending the two-list partition idea to multi-list partitions. At each generation step, the vocabulary set is divided into multiple lists. Based on the message to be inserted, the logits for tokens in a selected list are increased, while the token logits in all other lists remain unchanged.
@@
Instead of splitting the vocabulary set into different lists, Hu et al. (2024) introduce a method that randomly shuffles the order of all token probabilities within the interval $[0,1]$, setting the probabilities in the first half of the interval to 0 and doubling those in the second half. During the detection phase, a likelihood ratio test examines the significance of the likelihood that the given text is generated with the adjusted probability distribution. Wu et al. (2024) further generalizes this method by introducing a hyperparameter $\alpha\in[0,0.5]$, which controls the two cutoff points $\alpha$ and $1-\alpha$ within the interval $[0,1]$. The probability masses for the three resulting sub-intervals are adjusted accordingly.
@@
Watermarking by controlling sampling strategies. This category of watermarking methods inserts watermarks into the token sampling process by using watermark information to control the sampling of candidate tokens. For example, Christ et al. (2023) introduce a watermarking method that represents each token in the vocabulary set as a binary string of 0s and 1s. Next, a sequence of values from 0 to 1 is sampled uniformly. These values guide the token sampling process: if the predicted probability for a position in the binary string is larger than the corresponding pseudo-random value, that position is assigned a 1; otherwise, it is assigned a 0. Once all positions are determined, the token corresponding to the resulting binary string is sampled. Additionally, previous work (Kuditipudi et al., 2023) use a sequence of values randomly sampled from a uniform distribution between 0 and 1. The value controls the token sampling process through a decoder function, where the decoder function varies based on the sampling strategy. Hou et al. (2023) sample new sentences according to the original LLM until a sentence’s semantic value falls into the acceptance region. The acceptance region is predefined by randomly splitting the space of semantic embedding according to the context and the key.